# Integrated GARCH-GRU in Financial Volatility Forecasting

### v1

@@ 3.2.1 GARCH-GRU Model
The proposed integrated GARCH-GRU model builds upon the conventional GARCH(1,1) specification with constant mean assumption in Eq.1, which is formulated as:
@@
$\hat{h}_{t}$ $=\left(1-z_{t}\right)\odot h_{t-1}+z_{t}\odot\tilde{h}_{t}$ (19)
@@
The GARCH-GRU cell operates through dual processing streams. First, the GARCH component generates volatility estimates $g_{t}$ using recombination of past residuals and variances (Eq.15), where $\phi(\cdot)=W_{g}g_{t}+b_{g}$ denotes a linear projection aligning scalar GARCH output with GRU hidden state dimensions. Second, a modified GRU architecture processes sequential inputs through standard reset/update gates (Eqs.(16)-(19)), followed by additive integration of GARCH signals into the hidden state update (Eq.(20)). The learnable parameter $\gamma$ governs the contribution intensity of GARCH components, while element-wise addition enables granular modulation of individual hidden units. This design preserves GRU’s capacity to model nonlinear temporal dependencies while explicitly encoding volatility persistence patterns. A graphical representation of the integrated GARCH-GRU framework is shown in Figure.1.
@@ 3.2.2 GARCH-LSTM Model
To facilitate a systematic comparison between recurrent architectures in hybrid volatility modeling, we construct an integrated GARCH-LSTM model that parallels the design of the proposed GARCH-GRU framework, as formalized in Eqs.(21)–(28). Consistent with the GARCH-GRU’s integration strategy, the GARCH-LSTM model introduces a volatility-driven multiplicative gate derived from the standard GARCH(1,1) specification (Eq.(21)), which encodes key stylized facts of volatility, including clustering and persistence.
While both architectures incorporate GARCH signals to modulate neural dynamics, they diverge in the final fusion mechanism. In the GARCH-GRU model, the GARCH component is additively combined with the intermediate hidden state (Eq.(20)), allowing for element-wise modulation controlled by a learnable scaling parameter. In contrast, the GARCH-LSTM model adopts a multiplicative modulation strategy, partially inspired by the integration scheme proposed in Zhao et al. (2024). Specifically, the final hidden state $h_{t}$ is obtained by scaling the LSTM’s intermediate output $\tilde{h}_{t}$ with a learnable coupling parameter $w$ and a nonlinear transformation of the GARCH component $g_{t}$ (Eq.(28)). This multiplicative gating mechanism allows $w$ to adaptively calibrates the magnitude of GARCH contributions relative to the LSTM’s intrinsic temporal dynamics. As $w$ increases, the magnitude of $g_{t}$ becomes larger, thereby intensifying its modulation effect on the final hidden state and increasing its contribution to the model’s volatility forecast.
$g_{t}$ $=\omega_{0}+\alpha\epsilon_{t-1}^{2}+\beta\sigma_{t-1}^{2}$ (21)
@@ 4.1 Experimental Design
Datasets: We collect daily closing prices $p_{t}$ of three major U.S. stock indices, S&P 500, Dow Jones Industrial Average (DJI), and NASDAQ Composite, over the period from January 1, 2010, to December 31, 2019. This decade-long sample spans various market conditions, including periods of both low and high volatility. To construct the return series, we follow standard practices in the financial econometrics literature (Roh (2007), Bucci (2017), Kim and Won (2018), Bucci (2020)) and compute daily log returns as:
$r_{t}=log{\frac{p_{t}}{p_{t-1}}}$ (29)
@@
$\sigma_{t}$ $=\sqrt{\frac{1}{k}\sum_{i=0}^{k}(r_{t-i}-\mu_{k})^{2}}$ (30)
$\mu_{k}$ $=\frac{1}{k}\sum_{i=0}^{k}r_{t-i}$
The last 252 trading days, corresponding to roughly one calendar year, are used as the test set. Within the remaining training set, the final 20% is reserved for validation during model training and tuning.
Table 1: Statistics for Dataset
Dataset Count Mean Standard deviation Skewness Kurtosis ADF Non-zero Mean t-test LM
S&P 500 2414 0.0406 0.9307 -0.5746 4.5604 $-24.5061^{**}$ $2.1439^{*}$ $432.4055^{**}$
DJI 2414 0.0397 0.8858 -0.5697 3.9235 $-24.2261^{**}$ $2.2038^{*}$ $405.4800^{**}$
NASDAQ 2414 0.0501 1.0733 -0.5236 3.3275 $-24.0755^{**}$ $2.2947^{*}$ $342.0940^{**}$
‘*’ indicates significance at the 5% level, and ‘**’ indicates significance at the 1% level. ‘ADF’ represents the ADF test for unit root existence, ‘LM’ represents the ARCH Lagrange Multiplier test for heteroscedasticity existence.
Table.1 presents summary statistics and diagnostic tests for the return series of each index. All return series reject the null hypothesis of the Augmented Dickey-Fuller (ADF) test at the 1% significance level, confirming that the return series are stationary and do not contain unit roots. Additionally, the Lagrange Multiplier (LM) test for ARCH effects strongly rejects the null hypothesis of homoscedasticity for all three indices, indicating the presence of conditional heteroskedasticity, and thereby justifying the application of GARCH-type models. Furthermore, a two-sided t-test for non-zero mean is performed to assess the validity of the constant mean assumption adopted in volatility models. The test results indicate statistical significance at the 5% level across all datasets, thereby supporting the validity of the constant mean assumption adopted in our models for the subsequent empirical analysis.
@@
Model Configuration: The proposed integrated GARCH-GRU architecture is implemented by introducing custom modifications to the standard GRU cell, embedding conditional heteroskedasticity dynamics in a manner consistent with the GARCH(1,1) specification. A parallel implementation is applied to the integrated GARCH-LSTM model, serving as a comparative counterpart. This hybrid design enables the models to simultaneously capture nonlinear temporal dependencies and time-varying volatility structures inherent in financial time series.
To optimize model performance, hyperparameter tuning is conducted using the Optuna framework, leveraging a Bayesian optimization strategy for efficient exploration of the parameter space. To prevent overfitting, an early stopping criterion is applied based on validation set performance, with a patience threshold of 10 epochs. The mean squared error (MSE) is used as the monitoring metric. Model training is carried out using the Adam optimizer, which offers adaptive learning rates and robust convergence properties. This training configuration ensures both stability during optimization and strong generalization across varying forecasting horizons, supporting the empirical robustness of the proposed hybrid volatility models.
Evaluation Framework: The forecasting performance of the proposed model is evaluated using two complementary statistical accuracy metrics: Mean Squared Error (MSE) and Mean Absolute Error (MAE). These metrics are computed over the out-of-sample forecast horizon to quantify the deviation of predicted volatility from realized volatility. The Mean Squared Error is defined as
$\mathrm{MSE}:=\frac{1}{n}\sum_{t=1}^{n}\left(\hat{\sigma}_{t}-\sigma_{t}\right)^{2}$
where $\hat{\sigma}_{t}$ and $\sigma_{t}$ denote the predicted and realized volatility at time $t$, respectively. MSE penalizes larger forecast errors more heavily due to the form of the quadratic loss function, making it particularly sensitive to extreme deviations. In contrast, the Mean Absolute Error is calculated as
$\mathrm{MAE}:=\frac{1}{n}\sum_{t=1}^{n}\left|\hat{\sigma}_{t}-\sigma_{t}\right|$
which provides a more robust measure of average forecast accuracy that is less influenced by outliers. Together, MSE and MAE offer a comprehensive view of both the magnitude and variability of forecast errors, facilitating comparative evaluation across benchmark models and robustness checks.
In addition to these statistical loss metrics, we employ Value-at-Risk (VaR) forecasting as a complementary evaluation tool to assess the practical utility of the volatility predictions in risk management. VaR violations and violation ratios are analyzed to quantify the accuracy of the forecasted volatility in capturing downside risk. This dual focus on statistical precision and financial relevance enables a rigorous and well-rounded evaluation of the proposed hybrid model.
@@
Benchmarks: To rigorously evaluate the forecasting performance of the proposed integrated GARCH-GRU model, we compare it against four representative classes of benchmark models that are widely studied in the volatility forecasting literature. These benchmarks are selected to reflect a broad spectrum of modeling paradigms, encompassing classical econometric approaches, standard deep learning architectures, and both embedded and pipeline-based hybrid frameworks. The first benchmark class includes embedded hybrid models, specifically the architecture proposed in Zhao et al. (2024), which modifies the LSTM cell by replacing the output gate with a GARCH-type regression. This embedded integration enables the model to learn both volatility dynamics and sequential dependencies within a unified recurrent structure, facilitating end-to-end training. We refer to this model as ‘bl_GARCH_LSTM’, and designate it as the baseline hybrid benchmark, given its architectural similarity to our GARCH-GRU model and its innovative embedding of econometric dynamics within deep learning modules. The second benchmark class comprises pipeline hybrid models, where the output of a GARCH model, typically the estimated conditional variances, are used as input features for a recurrent neural network. This decoupled, two-stage integration strategy allows the GARCH component to capture short-term volatility clustering, while the LSTM component learns additional nonlinear and long-range dependencies. We denote this model as ‘pl_GARCH_LSTM’, representing a common and pragmatic approach in hybrid modeling that leverages the strengths of both statistical and machine learning techniques in a modular fashion. The third class consists of pure deep learning models, including standard GRU, LSTM, and Transformer architectures. These models are capable of learning rich temporal patterns and nonlinear dependencies, but do not incorporate any explicit econometric mechanisms for modeling volatility. As such, they serve as benchmarks for evaluating the added value of incorporating structured volatility dynamics into neural networks. The fourth and final class includes traditional econometric models, specifically GARCH(1,1) and GJR-GARCH(1,1,1). These models are well-established in the literature for capturing conditional heteroskedasticity and leverage effects in financial return series. Their inclusion provides a point of comparison for assessing the performance gains achieved through hybridization with neural networks. Together, these benchmark classes, comprising classical models, standard deep learning architectures, pipeline hybrids, and embedded hybrids, form a comprehensive evaluation framework. This allows us to position the proposed GARCH-GRU model within a broad methodological context, ensuring that observed performance improvements are both robust and interpretable, and reflect meaningful advances over existing approaches in the literature.
Architectural Validation: Prior to the construction of the integrated GARCH-GRU architecture, we conducted a rigorous validation of the custom neural network (NN) components to ensure their correctness and reliability. Since the hybrid model relies on manually implemented neural cells to allow GARCH integration at the architectural level, it was essential to verify that these custom components replicate the functional behavior of standard, well-tested NN implementations. This validation step was crucial to attribute any observed performance difference in the hybrid model solely to the intended GARCH-NN fusion, rather than unintended implementation artifacts. To assess implementation equivalence, we treated the custom and standard neural networks as two deterministic functions, $f_{1}(x)$ and $f_{2}(x)$, respectively. A set of randomly generated input vectors $x$ was passed through both networks, yielding output sequences $y_{1}=f_{1}(x)$ and $y_{2}=f_{2}(x)$. The output were then compared using a paired t-test, which evaluates whether the mean difference between paired observations from the two models is statistically significant. Formally, we tested the null hypothesis
$H_{0}:\mathbb{E}\left[y_{1}-y_{2}\right]=0$
against the alternative
$H_{1}:\mathbb{E}\left[y_{1}-y_{2}\right]\neq 0$
The results of the paired t-test are summarized in Table.2. For both GRU and LSTM architectures, the mean differences between the custom and standard implementations were effectively zero, and the corresponding p-values substantially exceeded conventional significance thresholds (e.g., $\alpha$ = 0.01). These results indicate that there is no statistically significant difference between the output of the custom and standard models.
Table 2: Paired t-test Comparing Custom and Standard GRU/LSTM Results
Model Mean Difference t-statistic p-value
GRU 0.000 $0.0195^{**}$ $0.9845^{**}$
LSTM -0.0003 $-0.7857^{**}$ $0.4339^{**}$
Significance tested at $\alpha=1\%$ and represented with ‘**’. Results indicate no statistically significant differences between custom and standard models.
These findings validate the correctness and reliability of the custom neural cell implementations. As such, the proposed GARCH-GRU model can be confidently attributed with structural improvements in forecasting performance, rather than artifacts introduced through computational discrepancies.

### v2

@@ 3.2.1 GARCH-GRU Model
The proposed GARCH-GRU model builds upon the conventional GARCH(1,1) specification with constant mean assumption in Eq.(1), which is formulated as:
@@
$\hat{h}_{t}$ $=\left(1-z_{t}\right)\odot\tilde{h}_{t}+z_{t}\odot h_{t-1}$ (19)
@@
The GARCH-GRU cell operates through dual processing streams. First, the GARCH component generates volatility estimates $g_{t}$ using recombination of past residuals and variances (Eq.15), where $\phi(u)=W_{g}u+b_{g}$ denotes a linear projection aligning scalar GARCH output with GRU hidden state dimensions. Second, a modified GRU architecture processes sequential inputs through standard reset/update gates (Eqs.(16)-(19)), followed by additive integration of GARCH signals into the hidden state update (Eq.(20)). The learnable parameter $\gamma$ governs the contribution intensity of GARCH components, while element-wise addition enables granular modulation of individual hidden units. The formulation in Eq.(20) can be equivalently expressed in its inverse tanh form as
$\frac{1}{\gamma}\tanh^{-1}\left(h_{t}\right)=g_{t}+\frac{1}{\gamma}\hat{h}_{t}$ (21)
This transformation reveals that the GARCH output $g_{t}$ and the recurrent component $\hat{h}_{t}$ contribute additively to the latent pre-activation of the hidden state. In this unsaturated representation, $g_{t}$ acts as a volatility-driven bias shift, while $\hat{h}_{t}$ captures feature-dependent temporal dynamics. The parameter $\gamma$ regulates their relative influence: larger values amplify the volatility contribution, whereas smaller values recover standard GRU dynamics. Because $\tanh^{-1}(\cdot)$ is monotone, this decomposition provides an interpretable mapping between econometric variance processes and neural hidden representations, clarifying how conditional volatility information modulates the recurrent state evolution. This design preserves GRU’s capacity to model nonlinear temporal dependencies while explicitly encoding volatility persistence patterns. A graphical representation of the integrated GARCH-GRU framework is shown in Figure.1.
@@ 3.2.2 GARCH-LSTM Model
To facilitate a systematic comparison between recurrent architectures in hybrid volatility modeling, we construct a GARCH-LSTM model that parallels the design of the proposed GARCH-GRU framework, as formalized in Eqs.(22)–(29). Consistent with the GARCH-GRU’s integration strategy, the GARCH-LSTM model introduces a volatility-driven multiplicative gate derived from the standard GARCH(1,1) specification (Eq.(22)), which encodes key stylized facts of volatility, including clustering and persistence.
While both architectures incorporate GARCH signals to modulate neural dynamics, they diverge in the final fusion mechanism. In the GARCH-GRU model, the GARCH component is additively combined with the intermediate hidden state (Eq.(20)), allowing for element-wise modulation controlled by a learnable scaling parameter. In contrast, the GARCH-LSTM model adopts a multiplicative modulation strategy, partially inspired by the integration scheme proposed in zhao2024garch. Specifically, the final hidden state $h_{t}$ is obtained by scaling the LSTM’s intermediate output $\tilde{h}_{t}$ with a learnable coupling parameter $w$ and a nonlinear transformation of the GARCH component $g_{t}$ (Eq.(29)). This multiplicative gating mechanism allows $w$ to adaptively calibrates the magnitude of GARCH contributions relative to the LSTM’s intrinsic temporal dynamics. As $w$ increases, the magnitude of $g_{t}$ becomes larger, thereby intensifying its modulation effect on the final hidden state and increasing its contribution to the model’s volatility forecast.
@@ 4.1 Experimental Design
Datasets: We collect daily closing prices $p_{t}$ of three major U.S. stock indices, S&P 500, Dow Jones Industrial Average (DJI), and NASDAQ Composite, over the period from January 1, 2010, to December 31, 2020. The sample covers multiple volatility regimes, from prolonged low-volatility intervals to sharp turbulence, including the 2020 COVID-19 crisis, providing a comprehensive testing ground for model evaluation. To construct the return series, we follow standard practices in the financial econometrics literature (roh2007forecasting; bucci2017forecasting; kim2018forecasting; bucci2020realized) and compute daily log returns as:
$r_{t}:=log{\frac{p_{t}}{p_{t-1}}}.$ (30)
@@
$\sigma_{t}$ $=\sqrt{\frac{1}{k}\sum_{i=1}^{k}(r_{t-i}-\mu_{k})^{2}}$ (31)
$\mu_{k}$ $=\frac{1}{k}\sum_{i=1}^{k}r_{t-i}$
The sample is divided chronologically to ensure strict out-of-sample evaluation. Observations prior to 2019 are used for model development, where rolling input windows are constructed from this pre-2019 segment to generate training sequences. And the last 20% of this pre-2019 data is reserved for validation during hyperparameter tuning. Out-of-sample evaluation is conducted primarily on the 2019 period, which represents a relatively stable, low-volatility market regime. To assess robustness in stressed environments, we additionally perform a complementary evaluation on the 2020 sample. The extended forecasts are generated using the same parameters calibrated from the pre-2019 setup, ensuring a genuine out-of-sample stress test reflective of the COVID-19 volatility shock.
Table 1: Summary Statistics of the Datasets
Dataset Count Mean Std. Dev. Skewness Kurtosis ADF Non-zero Mean t-test LM
S&P 500 2658 0.0422 1.1108 0.9122-0.9122 16.6378 $12.5060^{***}$ $1.9627^{**}$ $1042.7836^{***}$
DJI 2658 0.0382 1.1056 1.0128-1.0128 22.3001 $12.8978^{***}$ $1.7794^{*}$ $1059.0030^{***}$
NASDAQ 2658 0.0587 1.2327 0.8274-0.8274 11.1028 $14.5179^{***}$ $2.4537^{**}$ $887.2649^{***}$
  • ‘*’ indicates significance at the 10% level, ‘**’ at the 5% level, and ‘***’ at the 1% level. ‘ADF’ is the Augmented Dickey-Fuller test for unit-root existence. ‘LM’ is the ARCH Lagrange Multiplier test for heteroscedasticity.
Table.1 presents summary statistics and diagnostic tests for the return series of each index. All return series reject the null hypothesis of the Augmented Dickey-Fuller (ADF) test at the 1% significance level, confirming that the return series are stationary and do not contain unit roots. Additionally, the Lagrange Multiplier (LM) test for ARCH effects strongly rejects the null hypothesis of homoscedasticity for all three indices, indicating the presence of conditional heteroskedasticity, and thereby justifying the application of GARCH-type models. Furthermore, a two-sided t-test is conducted to assess whether the mean return is statistically different from zero. The resulting t statistics reject the null hypothesis of zero mean for all indices, supporting the constant but non-zero mean specification adopted in our volatility models for the subsequent empirical analysis.
@@
Model Configuration: The proposed GARCH–GRU and GARCH–LSTM architectures are constructed by introducing custom modifications to the standard GRU and LSTM cells, embedding conditional heteroskedasticity dynamics directly within their gating mechanisms in a manner consistent with the GARCH(1,1) specification. This integrated design allows each model to jointly learn nonlinear temporal dependencies and time-varying volatility structures in an end-to-end framework, thereby capturing both econometric stylized facts and complex recurrent dynamics inherent in financial time series. Model hyperparameters are tuned using the Optuna optimization framework, which employs a Bayesian search strategy to efficiently explore the parameter space. To mitigate overfitting, we apply early stopping based on validation performance, using a patience threshold of 20 epochs with mean squared error (MSE) as the monitoring criterion. Model training is performed using the Adam optimizer, providing adaptive learning rates and stable convergence properties. In addition, a dynamic learning-rate scheduler is employed, reducing the learning rate by a factor of two when validation performance stagnates. This training configuration ensures numerical stability and strong generalization across forecasting horizons, supporting the empirical robustness of the proposed GARCH-RNN hybrid volatility models.
Evaluation Framework:We assess the forecasting performance of the proposed models using a set of complementary statistical accuracy measures, including Mean Squared Error (MSE), Mean Absolute Error (MAE), out-of-sample $R^{2}$ and the Symmetric Mean Absolute Percentage Error (SMAPE). These metrics quantify the deviation between predicted volatility ($\hat{\sigma}_{t}$) from realized volatility ($\sigma_{t}$) over the out-of-sample horizon. MSE penalizes large errors more heavily and is therefore sensitive to extreme deviations, while MAE provides a more robust measure of average error magnitude. The out-of-sample $R^{2}$ evaluates improvement relative to a mean benchmark, offering a scale-adjusted indicator of predictive power. SMAPE captures proportional forecast errors, facilitating comparisons across volatility regimes that differ in scale.
To complement the statistical accuracy metrics, we evaluate the practical relevance of the volatility forecasts through Value-at-Risk (VaR) backtesting. In addition to examining VaR violations and violation ratios, we compute the pinball loss, which directly assesses the quality of quantile (VaR) forecasts by penalizing asymmetric under- and over-prediction. Incorporating the pinball loss provides a more sensitive assessment of tail-risk forecasting performance and strengthens the evaluation of the models’ utility in risk management applications.
@@
Benchmarks: To rigorously evaluate the forecasting performance of the proposed integrated GARCH-RNN framework, and in particular its GRU- and LSTM-based instantiations, we benchmark the models against several representative classes of approaches that span the main paradigms in the volatility forecasting literature. This set of benchmarks includes traditional econometric models, standard deep learning architectures, and alternative hybrid designs, thereby providing a comprehensive and interpretable basis for performance comparison. The first benchmark class consists of two-stage rolling hybrid models, where the GARCH and neural components are estimated sequentially rather than jointly. In this framework, denoted ‘bm_GARCH_LSTM’, the GARCH(1,1) model is used to generate adaptive volatility features, specifically $\alpha_{t}\varepsilon_{t}$ and $\beta_{t}\sigma_{t}$. Then these features are used as inputs to a LSTM network to forecast realized volatility over multiple horizons. This decoupled, pipeline-style integration enables the GARCH component to capture short-term volatility clustering and persistence, while the LSTM learns higher-order nonlinear temporal dependencies. The Two-Stage ‘bm_GARCH_LSTM’ thus serves as a representative and widely adopted hybrid benchmark, highlighting the difference between traditional pipeline integration and the fully embedded GARCH-RNN architectures proposed in this study. The second class consists of pure deep learning models, with particular emphasis on the Transformer architectures. These networks are capable of learning rich temporal patterns and nonlinear dependencies, but lack explicit econometric mechanisms for conditional heteroskedasticity. Their inclusion allows us to assess the incremental predictive value gained by embedding structured volatility dynamics directly within recurrent neural networks. The last class includes traditional econometric models, specifically GARCH(1,1) and GJR-GARCH(1,1,1). These models are well-established in the literature for capturing conditional heteroskedasticity and leverage effects in financial return series. Their inclusion provides a point of comparison for assessing the performance gains achieved through hybridization with neural networks. Together, these benchmark classes, comprising classical models, standard deep learning architectures, and pipeline-based hybrid designs, form a comprehensive evaluation framework. This allows us to position the proposed GARCH-RNN model within a broad methodological context, ensuring that observed performance improvements are both robust and interpretable, and reflect meaningful advances over existing approaches in the literature.
Prior to the construction of the integrated GARCH-GRU/LSTM architecture, we conducted a rigorous validation of the custom-implemented neural network architectures to ensure their correctness, reliability and equivalence to the standard Pytorch counterparts. The objective of these tests was to confirm that, when operating in a pure recurrent mode without any GARCH-specific components, the custom models yield outputs and gradients that are numerically identical to those produced by the canonical PyTorch implementations under the same initialization and inputs.
Both forward and backward equivalence checks were performed to establish full consistency. In the forward pass, each custom model and its PyTorch reference were provided with identical random input sequences and initial hidden states. The resulting outputs, including the complete hidden-state sequences and the final hidden (and cell) states, were compared under strict numerical tolerances, requiring relative and absolute errors below $10^{-6}$ to guarantee equality up to machine precision. In the backward pass, identical scalar losses were backpropagated through both networks, and the resulting gradients with respect to inputs, hidden states, and model parameters were compared element-wise, confirming that the gradient propagation was exactly matched. These tests were repeated under multiple random seeds to ensure reproducibility and robustness. All equivalence experiments passed within the prescribed numerical tolerances, demonstrating that our custom recurrent architectures (GRU and LSTM) are mathematically and computationally indistinguishable from the corresponding PyTorch implementations. This validation provides a solid foundation for the subsequent development of the GARCH-GRU and GARCH-LSTM hybrid models, ensuring that any observed differences in empirical performance can be attributed solely to the introduction of the GARCH mechanism rather than to inconsistencies in the underlying recurrent dynamics.